# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v2

@@ I Introduction
Despite its successful and widespread use, the Brody distribution has no theoretical foundation in the context of eigenvalue statistics (i.e., it is not derived from any physical or mathematical theory). Equation (3) is essentially a surmise, but because it is a surmise based on the well-established phenomenon of level repulsion, it is often described as being a “phenomenological” [36] or “semi-empirical” [27] distribution (“heuristic” [21] is another common descriptor). While the Brody distribution falls within the scope of random matrix theory (RMT) [11], its association with RMT is delusive since there are no preexisting random-matrix models whose eigenvalue spacing distributions are given by Eq. (3). In fact, the $2\times 2$ model of Nieminen and Muche [37] is (to the author’s knowledge) the only model that has been put forward as a candidate; unfortunately the spacing distribution for their model is not exactly the Brody distribution. Indeed, a long-standing criticism of the Brody distribution is that it does not derive from any known random-matrix model. The existence question [37] is resolved in the current paper, which presents a class of $2\times 2$ random-matrix models that has the Brody distribution as its exact eigenvalue spacing distribution. The uniqueness question is also addressed.
@@
The contents of the paper are as follows. In Sec. II, a general class of $2\times 2$ correlated random-matrix models is defined. The eigenvalue spacing distribution for this class of models is then derived in Sec. III with the result being the Brody distribution. Three simple sub-classes of models with varying matrix-element structures are defined and examined in Sec. IV; the significance of each sub-class is discussed therein and illustrative examples are provided. Additional examples involving complex matrices and eigenvalues are given in Sec. V. Some generalizations and extensions of the model introduced in Sec. II are considered in Sec. VI. Numerical results for selected real and complex examples from Secs. IV and V are then presented in Sec. VII. Comments (Sec. VIII), the Conclusion (Sec. IX), and several Appendices (A-E) complete the paper.
@@ II A General Class of $2\times 2$ Correlated Multivariate Models
where $m$ and $n$ are positive integers, $\{T,X,Y\}$ are independent random variables whose distributions will be specified below, $\{g_{1}(X),g_{2}(X)\}$ are functions of $X$ whose difference is a fixed (real or complex) number independent of $X$ (i.e., $g_{1}(X)-g_{2}(X)=\eta\in\mathbb{C}$, $\forall X$), $h(T)$ is any real function of $T$ having the property $h(T)\neq 0$ for all possible values of $T$, $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ are a constrained set of (real or complex) constants, and $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ ($p_{0}=0$) are a constrained set of real constants that can (but need not) have parametric dependencies. [^2] For example, some or all of $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ could depend on the parameter $q$. The distributions of $\{T,X,Y\}$ are as follows. Let the random variable $Y\sim\text{Exp}(\sigma_{\text{e}})$, that is, let $Y$ have an exponential distribution with scale parameter $\sigma_{\text{e}}>0$: [^3] In this paper, the common notation $f_{V}(v)$ will be used to refer to the p.d.f. of an arbitrary random variable $V$ with lower-case $v$ then denoting the density variable corresponding to $V$.
$f_{Y}(y;\sigma_{\text{e}})={1\over\sigma_{\text{e}}}\exp\left(-{y\over{\sigma_{\text{e}}}}\right),~{}y\geq 0.$ (7)
@@ III Derivation of the Eigenvalue Spacing Distribution
Utilizing the fact that the support of $Y$ is $[0,\infty)$ (note the restriction on the density variable $y$ in Eq. (7)), application of Eq. (10) to the model defined by Eqs. (6)-(8) yields:
@@
To determine the distribution of $S$, it is mathematically efficient to exploit the well-known functional relation between exponential and Weibull random variables (see, for example, Ref. [66]): if $Y\sim\text{Exp}(\sigma_{\text{e}})$, then the random variable $W\equiv Y^{1/\tau}\sim\text{Weibull}(\sigma=\sigma_{\text{e}},\tau)$. In words, this states that if $Y$ has an exponential distribution with scale parameter $\sigma_{\text{e}}>0$, then the random variable $W\equiv Y^{1/\tau}$, where $\tau>0$, has a Weibull distribution with scale parameter $\sigma=\sigma_{\text{e}}$ and shape parameter $\tau$. Thus,
$W\equiv Y^{1\over{q+1}}\sim\text{Weibull}\left(\sigma=\sigma_{\text{e}},\tau=q+1\right).$ (12)
Using (123) with parameter $\tau=q+1$, the p.d.f. of $S$ is thus:
$f_{S}(s)$ $=$ $f_{\sqrt{k}W}(s)={1\over\sqrt{k}}f_{W}\left(w={s\over\sqrt{k}};\sigma=\sigma_{\text{e}},\tau=q+1\right)$ (13)
@@
$\mu_{S}=\mu_{\sqrt{k}W}=\sqrt{k}\mu_{W}(\sigma=\sigma_{\text{e}},\tau=q+1)=\sqrt{k}\sigma_{\text{e}}^{1\over{q+1}}\Gamma\left({{q+2}\over{q+1}}\right).\quad$ (15)
@@ IV Three Sub-Classes of Models
where the constants $\{c_{j}:j=1,\ldots,6\}$ are generally different from zero. For the eigenvalues of $\mathcal{M}_{E}$ to have a Brody spacing distribution, various conditions on the constants $\{c_{j}\}$ and the power constants $\{a,b\}$ can be imposed such that condition (8) is satisfied. As examples, the random matrix
$\mathcal{M}_{E}^{(1)}=\left(\begin{array}[]{cc}e^{q}+c\log_{2}(|X|+1)&{1\over 2}k\left(1-Y^{1\over{q+1}}\right)\\ -{1\over 2}\left(1+Y^{1\over{q+1}}\right)&~{}e^{-q}+c\log_{2}\left({|X|+1\over 2}\right)\end{array}\right),$ (30)
where $k\equiv k(c,q)=(c+2\sinh(q))^{2}$, and the random matrix
@@
are two members of sub-class E. [^5] In the case of $\mathcal{M}_{E}^{(1)}$, the real constant $c$ is subject to the restriction $c\neq{-2\sinh(q)}$, but is otherwise arbitrary. Note that in each of $\mathcal{M}_{E}^{(1)}$ and $\mathcal{M}_{E}^{(2)}$, the off-diagonal matrix elements are independent of the diagonal matrix elements.
@@ IV.1.1 Illustrative Examples
A zero-trace subset of sub-class A exists that can be obtained by setting $X=0$ and fixing $c_{4}=-c_{1}$ ($c_{1}$ is then a free constant). For example, choosing
@@ IV.1.2 Significance of Sub-Class A
$\mathcal{M}^{(2)}_{I}=\left(\begin{array}[]{cc}X+Y^{\sin^{2}(q)\over{q+1}}&~{}Y^{\sin^{2}(q)-2\cos^{2}(q)\over{q+1}}-Y^{\sin^{2}(q)\over{q+1}}\\ -Y^{{\sin^{2}(q)+2\cos^{2}(q)\over{q+1}}}&X-Y^{\sin^{2}(q)\over{q+1}}\end{array}\right).$ (43)
@@ IV.1.4 A Related Class of Models
where the diagonal functions $\{d_{1}(X),d_{2}(X)\}$ are any two real functions of the random variable $X$ that satisfy $d_{1}(X)-d_{2}(X)=\eta\in\mathbb{R}$ ($\forall X$), the unequal constants $\{a,b\}$ satisfy condition (34e) (but are otherwise arbitrary), and the real constants $\{c_{1},c_{2},c_{3}\}$ satisfy the condition $c^{2}_{1}\eta^{2}+4c_{2}c_{3}>0$. Members of this class also satisfy condition (8) and hence have Brody spacing distributions.
@@ IV.2.1 Illustrative Examples
An interesting situation that can occur for each of Cases I-III is when one of the exponents is independent of $q$. In Cases I and III, this can result from simply prescribing $b=\pm(q+1)u$ and $a=\pm(q+1)v$ (respectively), where $u$ and $v$ are arbitrary $q$-independent constants or parameters. For Case II, this occurs when one of $\{a(q),b(q)\}$ is equal to ±$\pm(q+1)$; an example is:
$\mathcal{M}^{(2)}_{B(II)}=\left(\begin{array}[]{cc}-Y^{{1-q}\over{q+1}}&~{}-Y^{{1-q}\over{q+1}}+{1\over 4}Y\\ {1\over 4}Y^{{1-q}\over{q+1}}-Y&-Y\end{array}\right).$ (56l)
@@ IV.3 Sub-Class C
where the real constants $\{c_{j}:j=1,\ldots,6\}$ and power constants $\{a,b,c,d\}$ are subject to various sets of conditions such that condition (8) is satisfied. The constants are also subject to the following restrictions: (i) $\left\{c_{1},c_{3},c_{5},c_{6}\right\}$ are non-zero; and (ii) $\left\{a,b,c,d\right\}$ are non-zero and distinct.
@@ IV.3.1 Illustrative Examples
and the power constants $\{a,b,c,d\}$ satisfy the conditions
@@ IV.3.2 Significance of Sub-Class C
For any given order $n\geq 4$, there exist sub-classes of pruned models where all of the matrix elements are composed of strictly distinct random variables chosen from the set $\left\{Y^{{p_{1}}\over{q+1}},Y^{{p_{2}}\over{q+1}},\ldots,Y^{{p_{n}}\over{q+1}}:n\in\mathbb{N}\right\}$ (where the real parameters $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ are all non-zero and distinct). The significance of sub-class C is that it is the simplest such sub-class.
@@
where there are no restrictions on the constants $\{c_{j}:j=1,\ldots,8\}$ beyond those required to satisfy condition (8) but the power constants $\{a\neq 1,b,c,d\neq 1\}$ need still be distinct. [^9] With reference to model (6), both $\mathcal{M}_{C}$ and $\mathcal{M}_{G}$ are sub-classes of class $\mathcal{M}_{4}$. By virtue of their definitions, sub-classes A, B, and C, are disjoint. The restrictions on the power constants given here as part of the definition of sub-class G guarantee that sub-classes A and G are disjoint. As an example, one member of sub-class G is the random matrix:
@@ V.1 Complex Matrices with Real Eigenvalues
Condition (34d) for sub-class A and the conditions in (55a)-(55c) for sub-class B each constitute an underdetermined quadratic polynomial system. As previously discussed, these systems possess an infinite number of real and complex solutions. In defining sub-classes A, B, and C, all constants were stipulated to be real. This was done to simplify the conditioning of the constants, but is not a strict requirement. If the constants are permitted to be complex, then the defining conditions can be generalized so as to admit complex matrices with real eigenvalues. For example, condition (34d) for sub-class A would be replaced by the following three conditions:
@@ V.2 Complex-Conjugate Eigenvalues
Some illustrative examples having CC eigenvalues are given below (discriminant constants are calculated using the formulas given in Appendix B).
@@ VI.1 Generalization of the Exponents of $Y$
VI.2 Cross-Over Transitions Between Other RMT Universality Classes
@@
$f_{Y}(y;\sigma_{\text{g}},\xi=2)=\left({1\over\sigma_{\text{g}}^{2}}\right)y\exp\left(-{y\over\sigma_{\text{g}}}\right),~{}y\geq 0.$ (101)
Note that $Y$ has a “semi-Poisson” distribution (in the density variable $y$) when $\sigma_{\text{g}}=1/2$. It can be shown (see Appendix C) that when the distribution of $Y$ is given by Eq. (101) and condition (8) holds, then the (mean-scaled) spacing distribution for model (6) is:

### v3

@@ I Introduction
Despite its successful and widespread use, the Brody distribution has no theoretical foundation in the context of eigenvalue statistics (i.e., it is not derived from any physical or mathematical theory). Equation (3) is essentially a surmise, but because it is a surmise based on the well-established phenomenon of level repulsion, it is often described as being a “phenomenological” [36] or “semi-empirical” [27] distribution (“heuristic” [21] is another common descriptor). While the Brody distribution falls within the scope of random matrix theory (RMT) [11], its association with RMT has until now been groundless since there are no preexisting random-matrix models whose eigenvalue spacing distributions are given by Eq. (3). In fact, the $2\times 2$ model of Nieminen and Muche [37] is (to the author’s knowledge) the only model that has been put forward as a candidate; unfortunately the spacing distribution for their model is not exactly the Brody distribution. Indeed, a long-standing criticism of the Brody distribution is that it does not derive from any known random-matrix model. The existence question [37] is resolved in the current paper, which presents a class of $2\times 2$ random-matrix models that has the Brody distribution as its exact eigenvalue spacing distribution. The question of uniqueness is also addressed.
@@
The contents of the paper are as follows. In Sec. II, a general class of $2\times 2$ correlated random-matrix models is defined. The eigenvalue spacing distribution for this class of models is then derived in Sec. III with the result being the Brody distribution. Three simple sub-classes of models with varying matrix-element structures are defined and examined in Sec. IV; the significance of each sub-class is discussed therein and illustrative examples are provided. Additional examples involving complex matrices and eigenvalues are given in Sec. V. Some generalizations and extensions of the model introduced in Sec. II are considered in Sec. VI. Numerical results for selected real and complex examples from Secs. IV and V are then presented in Sec. VII. Comments (Sec. VIII), the Conclusion (Sec. IX), and several Appendices (A-D) complete the paper.
@@ II A General Class of $2\times 2$ Correlated Multivariate Models
where $m$ and $n$ are positive integers, $\{T,X,Y\}$ are independent random variables whose distributions will be specified below, $\{g_{1}(X),g_{2}(X)\}$ are functions of $X$ whose difference is a fixed (real or complex) number for all supported values of $X$ (i.e., $g_{1}(x)-g_{2}(x)=\eta\in\mathbb{C}$, $\forall x\in\text{supp}(X)$), $h(T)$ is any function of $T$ having the property $h(T)\neq 0$ for all supported values of $T$, $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ are a constrained set of (real or complex) constants, and $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ ($p_{0}=0$) are a constrained set of unequal real constants that can (but need not) have parametric dependencies. [^2] For example, some or all of $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ could depend on the parameter $q$. The distributions of $\{T,X,Y\}$ are as follows. Let the random variable $Y\sim\text{Exp}(\sigma_{\text{e}})$, that is, let $Y$ have an exponential distribution with scale parameter $\sigma_{\text{e}}>0$: [^3] In this paper, the common notation $f_{V}(v)$ will be used to refer to the p.d.f. of an arbitrary random variable $V$ with lower-case $v$ then denoting the density variable corresponding to $V$.
$f_{Y}(y;\sigma_{\text{e}})={1\over\sigma_{\text{e}}}\exp\left(-{y\over{\sigma_{\text{e}}}}\right),~{}y>0.$ (7)
@@ III Derivation of the Eigenvalue Spacing Distribution
Utilizing the fact that the support of $Y$ is $(0,\infty)$ (note the restriction on the density variable $y$ in Eq. (7)), application of Eq. (10) to the model defined by Eqs. (6)-(8) yields:
@@
To determine the distribution of $S$, it is mathematically efficient to exploit the well-known functional relation between exponential and Weibull random variables (see, for example, Ref. [66]): if $Y$ has an exponential distribution with scale parameter $\sigma_{\text{e}}>0$, then the random variable $W\equiv Y^{1/\tau}$, where $\tau>0$, has a Weibull distribution with scale parameter $\sigma_{\text{e}}$ and shape parameter $\tau$. Thus,
$W\equiv Y^{1\over{q+1}}\sim\text{Weibull}\left(\theta=\sigma_{\text{e}},\tau=q+1\right),$ (12)
where the general notation for the scale and shape parameters of the Weibull distribution given in Appendix A.1 has been utilized. Using (128) with parameter $\tau=q+1$, the p.d.f. of $S$ is thus:
$f_{S}(s)$ $=$ $f_{\sqrt{k}W}(s)={1\over\sqrt{k}}f_{W}\left(w={s\over\sqrt{k}};\theta=\sigma_{\text{e}},\tau=q+1\right)$ (13)
@@
$\mu_{S}=\mu_{\sqrt{k}W}=\sqrt{k}\mu_{W}(\theta=\sigma_{\text{e}},\tau=q+1)=\sqrt{k}\sigma_{\text{e}}^{1\over{q+1}}\Gamma\left({{q+2}\over{q+1}}\right).\quad$ (15)
@@ IV Three Sub-Classes of Models
where $a\neq b$, and the constants $\{c_{j}:j=1,\ldots,6\}$ are generally different from zero. For the eigenvalues of $\mathcal{M}_{E}$ to have a Brody spacing distribution, various conditions on the constants $\{c_{j}\}$ and the power constants $\{a,b\}$ can be imposed such that condition (8) is satisfied. For example, the random matrix
@@
is one member of sub-class E. Note that in the above example, the off-diagonal matrix elements are independent of the diagonal matrix elements.
@@ IV.1.1 Illustrative Examples
A cropped subset of sub-class A with vanishing trace can be obtained by setting $X=0$ and fixing $c_{4}=-c_{1}$ ($c_{1}$ is then a free constant). For example, choosing
@@ IV.1.2 Significance of Sub-Class A
$\mathcal{M}^{(2)}_{I}=\left(\begin{array}[]{cc}X+Y^{\sin^{2}\left(\phi\right)\over{q+1}}&~{}Y^{\sin^{2}\left(\phi\right)-2\cos^{2}\left(\phi\right)\over{q+1}}-Y^{\sin^{2}\left(\phi\right)\over{q+1}}\\[2.84544pt] -Y^{{\sin^{2}\left(\phi\right)+2\cos^{2}\left(\phi\right)\over{q+1}}}&X-Y^{\sin^{2}\left(\phi\right)\over{q+1}}\end{array}\right),$ (40)
where $\phi\equiv\phi(q)$ can be any real function of $q$ that is continuous and non-zero for all $q\in[0,1]$ (e.g., $\phi(q)=\pi^{q-1}$).
@@ IV.1.4 A Related Class of Models
where the diagonal functions $\{d_{1}(X),d_{2}(X)\}$ are any two real functions of the random variable $X$ that satisfy $d_{1}(X)-d_{2}(X)=\eta$, where $\eta$ is a real-valued constant, the unequal constants $\{a,b\}$ satisfy condition (31e) (but are otherwise arbitrary), and the real constants $\{c_{1},c_{2},c_{3}\}$ satisfy the condition $c^{2}_{1}\eta^{2}+4c_{2}c_{3}>0$. Members of this class also satisfy condition (8) and hence have Brody spacing distributions.
@@ IV.3 Sub-Class C
where the real constants $\{c_{j}:j=1,\ldots,6\}$ and power constants $\{a,b,c,d\}$ are subject to various sets of conditions such that condition (8) is satisfied. The constants are also subject to the following restrictions: (i) $\left\{c_{1},c_{3},c_{5},c_{6}\right\}$ are non-zero; and (ii) $\left\{a,b,c,d\right\}$ are distinct. [^8] Note that, by virtue of their definitions, sub-classes A, B, and C, are disjoint.
@@ IV.3.1 Illustrative Examples
and the unequal power constants $\{a,b,c,d\}$ satisfy the conditions
@@ IV.3.2 Significance of Sub-Class C
For any given order $n\geq 4$, there exist sub-classes of pruned models where all of the matrix elements are composed of strictly distinct random variables chosen from the set $\left\{Y^{{p_{1}}\over{q+1}},Y^{{p_{2}}\over{q+1}},\ldots,Y^{{p_{n}}\over{q+1}}:n\in\mathbb{N}\right\}$ (where the real parameters $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ are all non-zero and distinct). The significance of sub-class C is that it is the simplest such sub-class (under the auxiliary condition that $\left\{a,b,c,d\right\}$ are all non-zero).
@@
where there are no restrictions on the constants $\{c_{j}:j=1,\ldots,8\}$ beyond those required to satisfy condition (8) and the power constants $\{a\neq 1,b,c,d\neq 1\}$ are restricted from being equal. [^9] With reference to model (6), both $\mathcal{M}_{C}$ and $\mathcal{M}_{G}$ are sub-classes of class $\mathcal{M}_{4}$. The restrictions on the power constants given here as part of the definition of sub-class G guarantee that sub-classes A and G are disjoint. As an example, one member of sub-class G is the random matrix:
@@ V.1 Complex Matrices with Real Eigenvalues
Condition (31d) for sub-class A and the conditions in (52a)-(52c) for sub-class B each constitute an underdetermined quadratic polynomial system. As previously discussed, these systems possess an infinite number of real and complex solutions. In defining sub-classes A, B, and C, the constants $\{c_{j}\}$ were stipulated to be real. This was done to simplify the conditioning of the constants, but is not a strict requirement. If the constants $\{c_{j}\}$ are permitted to be complex, then the defining conditions can be generalized so as to admit complex matrices with real eigenvalues. For example, condition (31d) for sub-class A would be replaced by the following three conditions:
@@ V.2 Complex-Conjugate Eigenvalues
Some illustrative examples having CC eigenvalues are given below.
@@ VI.1.1 Illustrative Examples
Comment: $q$-Independent Exponents
For any sub-class of model (6) with $n\geq 2$, it is possible to prescribe $q$-dependent values for parameters $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ in such manner that at least one exponent (of $Y$) is independent of $q$. For example, in the case of sub-class C with conditions (66), prescribing the power constant $d=p(q+1)$, where $p\in\mathbb{R}\backslash\left[{1\over 3},{2\over 3}\right]$ is a real $q$-independent parameter, yields two $q$-independent exponents. As a specific example, choosing $p=2$ and constants $\{c_{j}\}$ as in model (72) yields:
$\mathcal{M}^{(4)}_{C}=\left(\begin{array}[]{cc}Y^{-{2q\over{q+1}}}&~{}-{1\over 2}Y^{4}\\ {1\over 2}\left(Y^{-{4(2q+1)\over{q+1}}}+1\right)&~{}-Y^{2}\end{array}\right).$ (100)
For another specific example, see model (82).
VI.2 Generalization of the Summand Coefficients $\{k_{m}\}$ in Model (6)
For given $n\in\mathbb{N}$ in model (6), the set $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ was defined to be a constrained set of (real or complex) constants. These coefficients are constrained in the sense that their values are restricted such that condition (8) is satisfied, but they are otherwise arbitrary. In particular, these coefficients need not be fixed numbers.
VI.2.1 Illustrative Example
As an illustrative example, consider the coefficients $\{c_{j}:j=1,\ldots,6\}$ of sub-class C, which in a previous example (c.f., Sec. IV.3.1) were stipulated to be real constants satisfying conditions (66a)-(66b). Setting
$c_{1}=\widetilde{c_{1}}p(V),~{}c_{4}={\widetilde{c_{4}}\over p^{2}(V)},~{}c_{5}=\widetilde{c_{5}}p^{2}(V),~{}c_{6}={\widetilde{c_{6}}\over p(V)}$ (101)
in (66a)-(66b) yields the transformed conditions
$\left(\widetilde{c_{1}}^{2}+4c_{3}\widetilde{c_{5}}\right)p^{2}(V)=0,~{}c_{2}=0,$ (102a)
${\left(\widetilde{c_{6}}^{2}+4c_{3}\widetilde{c_{4}}\right)\over p^{2}(V)}=0,~{}\widetilde{c_{1}}\widetilde{c_{6}}<0,$ (102b)
where the random variable $V$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{V}(v)$, and $p(V)$ is any real function of $V$ having the property $p(V)\neq 0$ for all supported values of $V$. For any given $p(V)$, system (102) again constitutes an underdetermined quadratic polynomial system possessing an infinite number of real (and complex) solutions. For example, one solution to system (102) is $\{\widetilde{c_{1}}=-1,c_{2}=0,c_{3}=-{1\over 2},\widetilde{c_{4}}=\widetilde{c_{5}}={1\over 2},\widetilde{c_{6}}=1\}$. Choosing power constants $\{a=0,b=4,c=-4,d=2\}$ (which satisfy conditions (66c)) and $p(V)=\cosh(qV)$ then yields:
$\mathcal{M}^{(5)}_{C}=\left(\begin{array}[]{cc}-\cosh(qV)&~{}-{1\over 2}Y^{4\over{q+1}}\\ {1\over 2}\left(\cosh^{2}(qV)Y^{-{4\over{q+1}}}+\text{sech}^{2}(qV)\right)&~{}\text{sech}(qV)Y^{2\over{q+1}}\end{array}\right).$ (105)
Note that in the above example the upper diagonal and upper off-diagonal matrix elements are mutually independent.
VI.3 Cross-Over Transitions Between Other RMT Universality Classes
@@
$f_{Y}(y;\sigma_{\text{g}},\xi=2)=\left({1\over\sigma_{\text{g}}^{2}}\right)y\exp\left(-{y\over\sigma_{\text{g}}}\right),~{}y>0.$ (106)
Note that $Y$ has a “semi-Poisson” distribution (in the density variable $y$) when $\sigma_{\text{g}}=1/2$. It can be shown (see Appendix B) that when the distribution of $Y$ is given by Eq. (106) and condition (8) holds, then the (mean-scaled) spacing distribution for model (6) is: